OpenStudy (anonymous):

1. Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. The water has absorbed the heat of the metal. So, qwater = qmetal Givens: Mass of metal (AI) = 27.776g Distilled water measurement = 26mL Distilled water temperature = 25.4 C Temperature of metal = 100.8 C Temperature of mixture (after metal is placed in calorimeter with distilled water) = 28.9 C Please and thank you!!

OpenStudy (anonymous):

2. Using the formula qmetal = m × c × ΔT, calculate the specific heat of the metal. Use the data from your experiment for the metal in your calculation.

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (anonymous):

My answer for question 1 is: Qwater = 27.776 * 4.18 J / (g * °C) * (25.4 * 38.9) = 6562.388 calories or 27475.410 Joules and for question 2 it's: Qmetal = 27.776 * c * (100.8 * 38.9) = -3.816 cal / g * C or -15.980 j / g * C

OpenStudy (bloomlocke367):

hmm, I'm not to good at chemistry. But @JFraser might be able to help.

OpenStudy (anonymous):

Ok, thanks :)

OpenStudy (bloomlocke367):

or @thomaster or @nincompoop.. try sending them a message and asking for help.

OpenStudy (anonymous):

Will do :P

OpenStudy (anonymous):

Can someone please check my work? @nincompoop @thomaster @Callisto @rvc

OpenStudy (anonymous):

gimme a sec I'll help you this was my favorite

OpenStudy (anonymous):

Yep you got it right :)

rvc (rvc):

Good job both of you :) All The Best

OpenStudy (jfraser):

\(\Delta T\) is the \(change\) in temperature, not the \(product\) of the two temperatures. You need to find the difference in temperature from the metal before and after, and the water before and after